# Detecting Out-of-Distribution Objects through Class-Conditioned Inpainting

### v3

@@ Abstract
Recent object detectors have achieved impressive accuracy in identifying objects seen during training. However, real-world deployment often introduces novel and unexpected objects, referred to as out-of-distribution (OOD) objects, posing significant challenges to model trustworthiness. Modern object detectors are typically overconfident, making it unreliable to use their predictions alone for OOD detection. To address this, we propose leveraging an auxiliary model as a complementary solution. Specifically, we utilize an off-the-shelf text-to-image generative model, such as Stable Diffusion, which is trained with objective functions distinct from those of discriminative object detectors. We hypothesize thatbu this fundamental difference enables the detection of OOD objects by measuring inconsistencies between the models. Concretely, for a given detected object bounding box and its predicted in-distribution class label, we perform class-conditioned inpainting on the image with the object removed. If the object is OOD, the inpainted image is likely to deviate significantly from the original, making the reconstruction error a robust indicator of OOD status. Extensive experiments demonstrate that our approach consistently surpasses existing zero-shot and non-zero-shot OOD detection methods, establishing a robust framework for enhancing object detection systems in dynamic environments.
Machine Learning, ICML
@@ 1 Introduction
Object detection systems have achieved remarkable progress in recent years, becoming integral to a wide range of applications in both online and offline settings. These include, but are not limited to, environmental monitoring (Beery et al., 2019), industrial manufacturing (Ahmad & Rahimi, 2022), and healthcare (Ragab et al., 2024). As these systems increasingly influence decision-making processes, ensuring their reliability and robustness has become a key challenge.
@@
One common failure mode arises when object detectors encounter object categories not seen during training, referred to as out-of-distribution (OOD) cases. Such occurrences are inevitable due to the dynamic and ever-changing nature of the real world. Unfortunately, modern object detectors often exhibit overconfidence, frequently misclassifying OOD objects as belonging to in-distribution (ID) categories, thereby resulting in false positives (Pathiraja et al., 2023). This overconfidence underscores a fundamental limitation: the internal responses of object detectors cannot reliably identify these errors. Consequently, alternative strategies are needed to enhance the trustworthiness of these systems.
@@
A popular strategy is to “rehearse” these error cases during detector training. For example, Munir et al. (2022); Pathiraja et al. (2023) introduced auxiliary losses to regularize model confidence, while Du et al. (2022b, a) proposed methods to learn more distinguishable detector features between ID and OOD categories. While promising, these strategies require retraining and are not widely implemented in pre-trained, publicly accessible object detectors. Given that many stakeholders rely on off-the-shelf detectors for their applications, there is a strong demand for effective post-hoc solutions to address OOD detection challenges.
@@
In this paper, we propose leveraging an additional pre-trained model to provide auxiliary information for effective OOD detection. Specifically, we utilize off-the-shelf text-conditioned generative models (Rombach et al., 2021; Saharia et al., 2022), which are trained to synthesize realistic images from text prompts, such as category names. We hypothesize that the distinct training objectives of object detectors and generative models lead to inconsistent responses for OOD cases. These inconsistencies thus serve as a robust post-hoc indicator for OOD detection, offering a practical and accessible solution to enhance the reliability of object detection systems without requiring retraining.
@@
More specifically, given a detected object bounding box and its predicted ID class label by the object detector, we propose performing class-conditioned inpainting for OOD detection. We mask the image portion within the box and use the ID class label as the prompt to resynthesize it. If the masked-out object is ID, the reconstructed image portion should be similar to the original, both visually and semantically. In contrast, for an OOD object, the inpainted portion prompted by the misclassified ID label will significantly deviate from the original. This discrepancy provides a valuable cue for distinguishing ID and OOD objects, requiring neither retraining nor access to ID or OOD training examples. We name our approach ZeRo-shot CONditional INpainting (RONIN), as illustrated in Figure 1.
@@
We evaluate RONIN on real-world benchmark datasets, including PASCAL VOC (Everingham et al., 2010), BDD-100K (Yu et al., 2020), MS-COCO (Lin et al., 2014), and OpenImages (Kuznetsova et al., 2020), covering a broad range of categorical domains and context scenarios, from indoor scenes to in-the-wild environments, for comprehensive evaluation of our proposed method. The empirical results demonstrate RONIN’s superior OOD detection performance compared to existing methods. Notably, RONIN excels in identifying near-OOD cases, which have semantic meanings closely related to some ID categories and are therefore particularly challenging to detect. This improvement is achieved by incorporating potential near-OOD labels as exclusions, resulting in larger reconstruction errors when the original object belongs to one of these near-OOD labels. Extensive ablation studies, analyses, and visualizations further validate the effectiveness of RONIN.
@@
  • Ease of implementation and compatibility: RONIN is straightforward to implement and compatible with most existing object detectors, including open-vocabulary object detectors (Liu et al., 2023a), which still rely on a pre-defined set of ID concepts to limit their scope.
@@
Remark.   While straightforward to implement, RONIN is far from a trivial contribution. It leverages the inconsistency between discriminative and generative models as an effective post-hoc cue for identifying OOD objects. Compared to Liu et al. (2023b), RONIN extends its applicability to object detection and eliminates the need to train a diffusion model on ID data. With the growing accessibility of pre-trained general-purpose models, utilizing them as auxiliary tools to enhance OOD detection for domain-specific models is both practical and logical.
One potential limitation of RONIN is its runtime. Diffusion models typically require longer runtimes compared to object detectors. However, this limitation can be mitigated by reducing the number of denoising steps. Since our goal is to detect OOD objects rather than generate high-quality images, a slight compromise in generation quality is acceptable. We examine this trade-off in  section 5 and show that reducing denoising steps significantly decreases runtime with minimal performance degradation.
Moreover, it is worth noting that object detectors are not always deployed in real-time settings. In many real-world scenarios, many real-time imaging often requires offline post processing such as camera-trap biology imaging in-the-wild, where performance and generalizability take precedence over runtime. By demonstrating consistently strong performances across outdoor and indoor scenarios through our experimental design, our framework RONIN effectively showcases its potential in real-time deployment.
Lastly, as generative modeling technologies continue to advance, we expect significant improvements in runtime efficiency, further enhancing the practicality of RONIN. By emphasizing flexibility, generability, and practicality both in our design choice and in our experimental performances, we believe RONIN approach showcases its valuable assets to contribute to the advancement of zero-shot out-of-distribution object detection.
@@ 2 Related Work
Out-of-distribution Detection.   The goal of the out-of-distribution (OOD) detection task is to determine whether a given sample belongs to a certain distribution. It has been widely studied at the image level, where a whole image is treated as a sample. Common approaches include leveraging classifier-specific information such as confidence scores (Hendrycks & Gimpel, 2016; Lee et al., 2017; Liang et al., 2018; DeVries & Taylor, 2018; Hsu et al., 2020; Liu et al., 2020; Wei et al., 2022) or learned features (Lee et al., 2018; Denouden et al., 2018; Tack et al., 2020; Sehwag et al., 2021; Xiao et al., 2021; Sun et al., 2022), or directly modeling an image distribution using generative models (Ren et al., 2019; Serrà et al., 2019; Xiao et al., 2020; Schlegl et al., 2017; Zong et al., 2018; Graham et al., 2023; Liu et al., 2023b; Li et al., 2023b). Recent works (Ming et al., 2022; Esmaeilpour et al., 2022; Wang et al., 2023) have also explored using CLIP (Radford et al., 2021) to identify OOD examples in a zero-shot manner, bypassing the need to learn from in-domain data explicitly.
OOD detection can also be extended to the object-level, where objects within an image are treated as individual samples. Most existing research in this setting focuses on training-time interventions. For instance, Du et al. (2022b, a) improve detector features to make them more separable between in-distribution (ID) and OOD data; Wilson et al. (2023) uses adversarial examples to train an MLP for classifying ID and OOD instances. In contrast, our approach RONIN explores the object-level setting through post-hoc interventions instead.
Text-to-Image Generative Models.   Recent advances in generative modeling have made large-scale text-to-image models widely available (Rombach et al., 2021; Ramesh et al., 2022; Saharia et al., 2022). These models exhibit a deep understanding of language and are highly effective at generating or editing high-quality images from diverse prompts, offering a promising approach to data synthesis across various tasks. Furthermore, recent works have shown that these models can also enhance discriminative tasks. For instance, Li et al. (2023a); Jaini et al. (2024) demonstrate that Stable Diffusion (Rombach et al., 2021) can function as an effective zero-shot classifier, achieving accuracy comparable to or surpassing CLIP and various trained discriminative classifiers. There has also been recent progress in applying text-conditioned diffusion models for image-level OOD detection (Du et al., 2023; Gao et al., 2023; Fang et al., 2024). Unlike these works, RONIN focuses on exploring the use of such models for object-level tasks.
@@ 3 Problem Formulation
We address the task of object-level out-of-distribution (OOD) detection, which is motivated by the fact that object detectors can recognize unseen objects as seen ones. Specifically, given an object detector trained to detect a pre-defined set of categories (e.g., different kinds of vehicles), we aim to identify the error cases when it wrongly recognizes a novel object (e.g., a wild animal) as one of the pre-defined-categories and detects it overconfidently. We refer to the pre-defined set of categories as the in-distribution (ID) classes, following (Du et al., 2022b); the novel categories as the OOD classes.
Formally, given an image $\bm{x}$ and an object detector $f(.;\ \theta)$ trained for the ID classes $\mathcal{Y}^{in}$, $f(\bm{x};\ \theta)$ outputs a list of bounding boxes $\bm{b}=\{b_{1},\ b_{2},\dots,\ b_{n}\}$ and their associated ID class labels $\bm{\hat{y}}=\{\hat{y}_{1}$, $\hat{y}_{2}$$\dots$, $\hat{y}_{n}$}, where $\hat{y}_{i}\in\mathcal{Y}^{in}$. Object-level OOD detection is then formulated as a binary classification problem, classifying whether the object within $b_{i}$ truly belongs to ID classes. Typically, one would develop a scoring function $\bm{g}$ such that given a bounding box and its predicted label $(b,\ \hat{y})$, $\bm{g}$ gives a higher score $\bm{g}(b,\ \hat{y})$ if $b$ outlines an ID object and a lower score if it outlines an OOD object.
Existing works (Du et al., 2022b, a) proposed a specific training process for the object detector so that its responses (e.g., the feature vector of each bounding box) would better distinguish between ID and OOD classes. However, doing so requires re-training the object detector, assuming access to the original training data and making it infeasible to off-the-shelf object detectors.
Aim.   To address this limitation, we aim to design an OOD detection mechanism that is post-hoc, without the need to modify (e.g., fine-tune) the pre-trained object detector, and zero-shot (Ming et al., 2022), without the need to access the original training data or any ID-class data. This sharply contrasts several prior post-hoc methods that need ID-class data (Ren et al., 2019; Xiao et al., 2020).
Approach.   The emergence of vision-language foundation models (Radford et al., 2021; Li et al., 2022; Liu et al., 2024a) trained on abundant images and free-form text pairs has gradually removed the boundary between the closed-set and open-set settings Li et al. (2023a). For example, while not perfect, CLIP (Radford et al., 2021) can match an image with unbounded concepts. Such a zero-shot capability has been leveraged in prior work to detect OOD samples given a set of ID concepts (Ming et al., 2022; Esmaeilpour et al., 2022; Wang et al., 2023). In this work, we leverage another kind of foundation model, text-to-image generative models (Rombach et al., 2021), capable of generating images given free-form text. While not designed for object detection nor specifically optimized for the ID data, we surmise their built-in, generic capability would facilitate object-level OOD detection.
@@ 4 RONIN: Zero-Shot OOD Conditional Inpainting
Figure 2: Overall framework of RONIN with two main components: (i) Class-conditioned Inpainting model and (ii) Vision-language Triplet Measurement. Given an input image with bounding boxes and in-distribution class labels, RONIN masks the objects within each bounding box and conditions the inpainting model to reconstruct them based on their predicted labels. The inpainted outcome is then compared to the original using the vision-language triplet measurement to evaluate semantic and visual alignment, enabling zero-shot OOD detection.
4.1 Theoretical motivation
Our proposed framework, RONIN, is developed based on the assumption that when encountering unseen OOD objects, generative and object detection have different options due to different training objectives. Generally, object detections are trained to separate foreground objects from the background and classify them into one of the pre-defined ID categories—i.e., learning $P(y|x)$. By minimizing this objective, they learn soft decision boundaries between categories, which can overconfidently classify unseen foreground objects into one seen label during inference. In contrast, generally trained generative models aim to model the underlying data distribution $P(x,y)$ per category by capturing their nuanced characteristics. When a sample deviates from the training distribution it learns, it can facilitate identifying this as an outlier. Based on these distinct objectives, we leverage generative models as an auxiliary indicator over object detection outcomes to enable effective zero-shot out-of-distribution detection.
We present the general framework of RONIN in Figure 2. Given an input object with its predicted bounding box and ID label, RONIN utilizes off-the-shelf class-conditioned inpainting models to regenerate the object, conditioning the process by the label to align it with the in-distribution domain. The inpainted outcome is then compared to the original object under a vision-language triplet similarity measurement. High similarity scores reflect strong visual and semantic alignment, suggesting the object is in-distribution, whereas lower scores indicate significant deviation, indicating out-of-distribution. Below, we provide a detailed component of our approach.
@@ 4.2 Class-conditional Inpainting
Our key idea behind RONIN is to harness diffusion models for inpainting, refine predicted objects, and bring them closer to the ID domain. Specifically, given a detected bounding box $\bm{b}$ and a predicted ID class $\hat{y}\in\mathcal{Y}^{in}$ by the object detector, we generate a masked region $\bm{m}$ within $\bm{b}$ and use $\hat{y}$ as a condition for reconstruction. Since $\bm{m}$ is always contained within $\bm{b}$, the inpainting model receives sufficient context from both the background and the detected object as hints for reconstructing the original object. This ensures that the inpainted outcomes closely align with the original ID objects, not only in semantic meaning but also in visual appearance, while still distinct in OOD cases.
More specifically, let the semantic background $\bm{s}$ be the entire image region excluding the masked region $\bm{m}$, RONIN takes input of $\bm{m}$ (also known as the inpainting mask), the predicted label $\bm{c}=\hat{y}$, and the semantic background $\bm{s}$ for inpainting. We choose to use pre-trained text-conditioned diffusion models such as DDPM (Ho et al., 2020). The diffusion models starts with a random Gaussian noise over $\bm{m}$ and perform the reverse diffusion process with $T$ steps, guided by $\bm{c}$, to obtain the outcome image $\bm{x}_{inp}={\bm{x}}_{0}$, as in Equation 1, where $\mu_{\theta}$ and $\Sigma_{\theta}$ are priorly learned by pre-training a neural network with parameter $\theta$.
@@
While the inpainted objects are synthesized based on $\hat{y}$ over the foreground region $\bm{m}$, the information retained from $\bm{s}$ guides the DDPM to preserve some visual features of the original object, as described in Equation 2. As a result, the inpainted outcome object retains the same visual and semantic characteristics as the original object, despite being presented differently. This enables RONIN to generate objects closely aligned with the originals in ID cases and significantly deviating in OOD cases, as shown in Figure 2. We further analyze the correlation between the bounding box $\bm{b}$ and the inpainted mask $\bm{m}$ under visual and semantic alignment perspectives in Appendix C.
Class-wise Masking and Inpainting.   Intuitively, a natural approach to inpainting is treating each object independently, or “object-wise” inpainting. However, this is highly inefficient when many objects share the same predicted label and denoising condition, making the process redundant. To address this, we propose “class-wise” inpainting, where objects are grouped by the predicted label $\hat{y}$ for a single inpainting process, ensuring the computational cost at most scales with $\mathcal{Y}^{in}$ in the worst case. This significantly reduces computational cost while maintaining performance, as the shared denoising condition ensures effective synthesis. We further evaluate the performance between the two strategies in our ablation study.

### v4

@@ Abstract
Recent object detectors have achieved impressive accuracy in identifying objects seen during training. However, real-world deployment often introduces novel and unexpected objects, referred to as out-of-distribution (OOD) objects, posing significant challenges to model trustworthiness. Modern object detectors are typically overconfident, making it unreliable to use their predictions alone for OOD detection. To address this, we propose leveraging an auxiliary model as a complementary solution. Specifically, we utilize an off-the-shelf text-to-image generative model, such as Stable Diffusion, which is trained with objective functions distinct from those of discriminative object detectors. We hypothesize that this fundamental difference enables the detection of OOD objects by measuring inconsistencies between the models. Concretely, for a given detected object bounding box and its predicted in-distribution class label, we perform class-conditioned inpainting on the image with the object removed. If the object is OOD, the inpainted image is likely to deviate significantly from the original, making the reconstruction error a robust indicator of OOD status. Extensive experiments demonstrate that our approach consistently surpasses existing zero-shot and non-zero-shot OOD detection methods, establishing a robust framework for enhancing object detection systems in dynamic environments. Our implementation is available at https://github.com/quanghuy0497/RONIN.
@@ 1 Introduction
Object detection systems have achieved remarkable progress in recent years, becoming integral to a wide range of applications in both online and offline settings. These include, but are not limited to, environmental monitoring [beery2019efficient], industrial manufacturing [ahmad2022deep], and healthcare [Ragab2024ACS]. As these systems increasingly influence decision-making processes, ensuring their reliability and robustness has become a key challenge.
@@
One common failure mode arises when object detectors encounter object categories not seen during training, referred to as out-of-distribution (OOD) cases. Such occurrences are inevitable due to the dynamic and ever-changing nature of the real world. Unfortunately, modern object detectors often exhibit overconfidence, frequently misclassifying OOD objects as belonging to in-distribution (ID) categories, thereby resulting in false positives [pathiraja2023multicalib]. This overconfidence underscores a fundamental limitation: the internal responses of object detectors cannot reliably identify these errors. Consequently, alternative strategies are needed to enhance the trustworthiness of these systems.
@@
A popular strategy is to “rehearse” these error cases during detector training. For example, [munir2022towards, pathiraja2023multicalib] introduced auxiliary losses to regularize model confidence, while [du2022vos, du2022siren] proposed methods to learn more distinguishable detector features between ID and OOD categories. While promising, these strategies require retraining and are not widely implemented in pre-trained, publicly accessible object detectors. Given that many stakeholders rely on off-the-shelf detectors for their applications, there is a strong demand for effective post-hoc solutions to address OOD detection challenges.
@@
In this paper, we propose leveraging an additional pre-trained model to provide auxiliary information for effective OOD detection. Specifically, we utilize off-the-shelf text-conditioned generative models [rombach2021highresolution, saharia2022imagen], which are trained to synthesize realistic images from text prompts, such as category names. We hypothesize that the distinct training objectives of object detectors and generative models lead to inconsistent responses for OOD cases. These inconsistencies thus serve as a robust post-hoc indicator for OOD detection, offering a practical and accessible solution to enhance the reliability of object detection systems without requiring retraining.
@@
More specifically, given a detected object bounding box and its predicted ID class label by the object detector, we propose performing class-conditioned inpainting for OOD detection. We mask the image portion within the box and use the ID class label as the prompt to resynthesize it. If the masked-out object is ID, the reconstructed image portion should be similar to the original, both visually and semantically. In contrast, for an OOD object, the inpainted portion prompted by the misclassified ID label will significantly deviate from the original. This discrepancy provides a valuable cue for distinguishing ID and OOD objects, requiring neither retraining nor access to ID or OOD training examples. We name our approach ZeRo-shot CONditional INpainting (RONIN), as illustrated in Fig. 1.
@@
We evaluate RONIN on real-world benchmark datasets, including PASCAL VOC [everingham2010pascal], BDD-100K [yu2020bdd100k], MS-COCO [lin2014microsoft], and OpenImages [kuznetsova2020open], covering indoor, outdoor, and in-the-wild scenarios for comprehensive evaluation. The empirical results demonstrate RONIN’s superior OOD detection performance compared to existing methods. Notably, RONIN shows strong potential in near-OOD scenarios by incorporating potential nearest non-ID concepts as exclusions, which leads to larger reconstruction errors when the original object belongs to one of these concepts and thus address near-OOD detection. Extensive ablation studies, analyses, and visualizations further validate the effectiveness of RONIN.
@@
  • Ease of implementation and compatibility: RONIN is straightforward to implement and compatible with most existing object detectors, including open-vocabulary object detectors [liu2023grounding], which still rely on a pre-defined set of ID concepts to limit their scope.
@@
Remark.   While straightforward to implement, RONIN is far from a trivial contribution. It effectively leverages the inconsistency between discriminative and generative models as a post-hoc signal for identifying OOD objects. Unlike [liu2023lmd], RONIN extends applicability to object detection without requiring diffusion models to be trained on ID data. With the growing availability of powerful pre-trained models, using them as auxiliary tools to enhance OOD detection in domain-specific pipelines is both practical and scalable.
A potential limitation of RONIN lies in runtime, as diffusion models are generally slower than object detectors. However, this can be mitigated by reducing denoising steps—since our focus is on detecting OOD objects, not generating high-fidelity images, a slight quality trade-off is acceptable. We investigate this in Sec. 5, showing that fewer denoising steps significantly reduce runtime with minimal performance drop.
Importantly, object detectors are not always used in real-time. Many applications, such as wildlife monitoring with camera traps, involve offline post-processing where accuracy and generalizability matter more than speed. Our experiments across indoor and outdoor scenes validate RONIN ’s strong performance in such settings, highlighting its readiness for practical deployment.
As generative modeling continues to evolve—with efficiency gains from methods like one-step diffusion—RONIN stands to benefit further. By prioritizing flexibility, generalizability, and practicality in both design and empirical results, RONIN offers a meaningful contribution to zero-shot out-of-distribution object detection.
@@ 2 Related Work
Out-of-distribution Detection.   The goal of the out-of-distribution (OOD) detection task is to determine whether a given sample belongs to a certain distribution. It has been widely studied at the image level, where a whole image is treated as a sample. Common approaches include leveraging classifier-specific information such as confidence scores [hendrycks2016baseline, lee2017training, liang2018odin, devries2018learning, hsu2020generalizedodin, liu2020energy, wei2022logitnorm] or learned features [lee2018simple, denouden2018improving, tack2020csi, sehwag2021ssd, xiao2021we, Sun2022knn], or directly modeling an image distribution using generative models [ren2019likelihood, serra2019input, xiao2020likelihoodregret, schlegl2017unsupervised, zong2018deep, Graham2023denoising, liu2023lmd, li2023rethinking]. Recent works [ming2022delving, esmaeilpour2022zero, wang2023clipn] have also explored using CLIP [radford2021clip] to identify OOD examples in a zero-shot manner, bypassing the need to learn from in-domain data explicitly.
OOD detection can also be extended to the object-level, where objects within an image are treated as individual samples. Most existing research in this setting focuses on training-time interventions. For instance, [du2022vos, du2022siren] improve detector features to make them more separable between in-distribution (ID) and OOD data; [wilson2023safe] uses adversarial examples to train an MLP for classifying ID and OOD instances. In contrast, our approach RONIN explores the object-level setting through post-hoc interventions instead.
Text-to-Image Generative Models.   Recent advances in generative modeling have made large-scale text-to-image models widely available [rombach2021highresolution, ramesh2022dalle, saharia2022imagen]. These models exhibit a deep understanding of language and are highly effective at generating or editing high-quality images from diverse prompts, offering a promising approach to data synthesis across various tasks. Furthermore, recent works have shown that these models can also enhance discriminative tasks. For instance, [li2023your, jaini2023intriguing] demonstrate that Stable Diffusion [rombach2021highresolution] can function as an effective zero-shot classifier, achieving accuracy comparable to or surpassing CLIP and various trained discriminative classifiers. There has also been recent progress in applying text-conditioned diffusion models for image-level OOD detection [du2023dream, gao2023diffguard, fang2024ddos]. Unlike these works, RONIN focuses on exploring the use of such models for object-level tasks.
@@ 3 Problem Formulation
We address the task of object-level out-of-distribution (OOD) detection, which is motivated by the fact that object detectors can recognize unseen objects as seen ones. Specifically, given an object detector trained to detect a predefined set of categories (e.g. different kinds of vehicles), we aim to identify the error cases when it wrongly recognizes a novel object (e.g. a wild animal) as one of the predefined categories and detects it overconfidently. We refer to the pre-defined set of categories as the in-distribution (ID) classes, following [du2022vos]; the novel categories as the OOD classes.
[…]
Existing works [du2022vos, du2022siren] proposed a specific training process for the object detector so that its responses (e.g. the feature vector of each bounding box) would better distinguish between ID and OOD classes. However, doing so requires re-training the object detector, assuming access to the original training data and making it infeasible to off-the-shelf object detectors.
Aim.   To address this limitation, we aim to design an OOD detection mechanism that is post-hoc, without the need to modify (e.g. fine-tune) the pre-trained object detector, and zero-shot [ming2022delving], without the need to access the original training data or any ID-class data. This sharply contrasts several prior post-hoc methods that need ID-class data [ren2019likelihood, xiao2020likelihoodregret].
Approach.   The emergence of vision-language foundation models [radford2021clip, li2022blip, liu2024visual] trained on abundant images and free-form text pairs has gradually removed the boundary between the closed-set and open-set settings [li2023your]. For example, while not perfect, CLIP [radford2021clip] can match an image with unbounded concepts. Such a zero-shot capability has been leveraged in prior work to detect OOD samples given a set of ID concepts [ming2022delving, esmaeilpour2022zero, wang2023clipn]. In this work, we leverage another kind of foundation model, text-to-image generative models [rombach2021highresolution], capable of generating images given free-form text. While not designed for object detection nor specifically optimized for the ID data, we surmise their built-in, generic capability would facilitate object-level OOD detection.
@@ 4 RONIN: Zero-Shot OOD Conditional Inpainting
Figure 2: Overall framework of RONIN, including (i) Class-conditioned Inpainting and (ii) Vision-language Triplet Measurement. Given an image with bounding boxes and predicted labels, RONIN masks and reconstructs objects via inpainting, then evaluates alignment using triplet similarity for zero-shot OOD detection.
Our proposed framework, RONIN, builds on the assumption that generative models and object detectors behave differently when encountering unseen OOD objects due to their distinct training objectives. Object detectors are trained to separate foreground from background and classify objects into pre-defined ID categories, learning $P(y|x)$. This results in soft decision boundaries that can overconfidently assign unseen objects to seen classes. In contrast, generative models aim to capture the full data distribution $P(x,y)$ by modeling nuanced category characteristics, making them sensitive to deviations from training data and thus effective at flagging outliers. Leveraging this, we use generative models as auxiliary signals to improve zero-shot OOD detection in conjunction with object detection outputs.
RONIN is illustrated in Fig. 2. Given an object with its predicted bounding box and ID label, RONIN uses off-the-shelf class-conditioned inpainting models to regenerate the object, aligning it with the in-distribution domain. The regenerated image is then compared with the original using a vision-language triplet similarity metric. High similarity implies visual and semantic consistency with the ID category (in-distribution), while low similarity suggests a mismatch (out-of-distribution). We detail each component of this approach below.
@@ 4.1 Class-conditional Inpainting
Our key idea behind RONIN is to harness diffusion models for inpainting, refine predicted objects, and bring them closer to the ID domain. Specifically, given a detected bounding box $\bm{b}$ and a predicted ID class $\hat{y}\in\mathcal{Y}^{in}$ by the object detector, we generate a masked region $\bm{m}$ within $\bm{b}$ and use $\hat{y}$ as a condition for reconstruction. Since $\bm{m}$ is always contained within $\bm{b}$, the inpainting model receives sufficient context from both the background and the detected object as hints for reconstructing the original object. This ensures that the inpainted outcomes closely align with the original ID objects, not only in semantic meaning but also in visual appearance, while still being distinct in OOD cases.
More specifically, let the semantic background $\bm{s}$ be the entire image region excluding the masked region $\bm{m}$, RONIN takes input of $\bm{m}$ (also known as the inpainting mask), the predicted label $\bm{c}=\hat{y}$, and the semantic background $\bm{s}$ for inpainting. We choose to use pre-trained text-conditioned diffusion models such as DDPM [ho2020denoising]. The diffusion models starts with a random Gaussian noise over $\bm{m}$ and perform the reverse diffusion process with $T$ steps, guided by $\bm{c}$, to obtain the outcome image $\bm{x}_{inp}={\bm{x}}_{0}$, as in Eq. 1, where $\mu_{\theta}$ and $\Sigma_{\theta}$ are priorly learned by pre-training a neural network with parameter $\theta$.
@@
While the inpainted objects are synthesized based on $\hat{y}$ over the foreground region $\bm{m}$, the information retained from $\bm{s}$ guides the DDPM to preserve some visual features of the original object, as described in Eq. 2. As a result, the inpainted outcome object retains the same visual and semantic characteristics as the original object, despite being presented differently. This enables RONIN the generation of objects closely aligned with the originals in ID cases and significantly deviating in OOD cases, as shown in Fig. 2. We further analyze how varying the inpainted mask $\bm{m}$ based on $\bm{b}$ impacts OOD detection in Sec. 6.
Class-wise Masking and Inpainting.   Intuitively, a natural approach to inpainting is treating each object independently, or “object-wise” inpainting. However, this is highly inefficient when many objects share the same predicted label and denoising condition, making the process redundant. To address this, we propose “class-wise” inpainting, where objects are grouped by the predicted label $\hat{y}$ for a single inpainting process, ensuring the computational cost at most scales with $\mathcal{Y}^{in}$ the worst case. This significantly reduces computational cost while maintaining performance, as the shared denoising condition ensures effective synthesis. We further evaluate the two strategies in Sec. 6.